# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ III-B3 Significance-oriented modulation policy design
Modulation techniques is crucial in the wireless communication system, for they can enhance the reliability, spectral efficiency, and data transmission rate. However, the modulation technique design encounters challenges in efficiently utilizing limited spectrum resources to ensure timeliness in the SAGSIN. To address this, NOMA has emerged as a promising solution, which enables multiple users to share time-frequency resources by using power domain multiplexing. By adopting power allocation (i.e., assigning varying power levels to different users) which is a modulation policy, NOMA facilitates simultaneous message transmission of multiple users. Allocating more power to transmitted data of higher significance not only ensures reliable transmission of significant data but also enhances the timeliness of the system by directly influencing data rates and system capacity. Various techniques are employed to enhance the timeliness performance of NOMA systems. [65] considers the problem of energy-efficient scheduling for minimizing the AoI in an opportunistic NOMA/orthogonal-multiple-access (NOMA/OMA) downlink broadcast wireless network. The initial step involves formulating a resource allocation problem to minimize the average AoI in the network. To address energy-efficiency considerations, the work takes into account both a long-term average power constraint and a maximum power constraint. Then the Lyapunov framework is employed to approximate the original problem as a queue stability problem. The obtained single time-slot optimization problem is further decomposed into power allocation and user scheduling subproblems. An efficient piece-wise linear approximation is utilized to solve the non-convex power allocation subproblem. Unlike the approach discussed in [65] that focuses solely on NOMA and OMA transmission schemes, [66] proposes an adaptive NOMA/OMA/cooperative-SWIPT-NOMA transmission scheme. In the NOMA scheme, the base station (BS) is responsible for allocating power to each user. However, in the cooperative-SWIPT-NOMA scheme, the BS not only allocates power to each user but also determines the power splitting coefficient, which determines the amount of harvested energy. The authors focus on a short packet communication system within a wireless network, where a BS transmits timely status updates to the users. The objective is to minimize the weighted average AoI by judiciously selecting appropriate multiple access techniques and corresponding power allocations. The resource allocation problem for the adaptive transmission scheme is formulated as an MDP and solved by iteration algorithms. Similar to [66], the authors in [67] investigate the AoI performance in a downlink wireless communication system, while HARQ-CC aided NOMA technology is employed in [67]. Furthermore, [67] considers a more flexible scenario where both power allocation decisions and retransmission decisions are made adaptively to minimize the system average AoI in each time slot. The authors notice the unfairness between users causes by the average-AoI-optimal policy. To address this issue, they formulate a problem to minimize the user’s maximum expected AoI, ensuring fairness by preventing the farther user with weak channel gains from being deprived of timely service. It can be proven that a stationary optimal policy exists for the transformed MDP problem. The aforementioned works primarily focus on ground-based wireless communication scenarios, while [68, 69] consider the satellite-based IoT (S-IoT) network. In particular, the authors in [68] propose an AoI-minimal resource allocation scheme in NOMA-based S-IoT downlink network. In this system, the satellite sends timely status updates to multiple user equipment (UE) devices. To ensure the freshness of the status updates in the network, an AoI optimization problem is formulated, taking into account long-term average power, peak power, and throughput constraints. The problem is then transformed into a series of online power allocation problems using Lyapunov optimization tools. Considering the limited computing resources of satellites, the particle swarm optimization (PSO) algorithm is employed to solve the non-convex optimization problem with linear computational complexity, named NOMA-AoI scheme. In the PSO algorithm, each particle means a possible solution which is the power allocation factors. By iteratively updating the velocities and positions of particles based on their historical information and the globally best information, the PSO algorithm explores the search space to find the optimal power allocation strategy. Based on [68], the authors in [69] further consider the network stability constraint in AoI optimization problem under NOMA-based S-IoT network. The problem is then transformed into three queue stability problems using the Lyapunov framework, which also converts the optimization problem into a series of single time slot deterministic optimization problems. The weights for the queue backlog and channel conditions are derived using the ListNet algorithm, a machine-learning-based approach, to obtain an optimized power allocation order with linear complexity. Similar to [68], the proposed NOMA long-term AoI minimization power allocation problem is also addressed using the PSO algorithm.
@@ III-C Lessons Learned from This Section
This section focuses on significance-oriented sampling, coding, and modulation policies to release heavy burden caused by massive source data and ensure that only significant data are sampled and transmitted for further processing and actuation. In fact, in the SAGSIN scenarios, although data generated by distributed edge devices are massive and multi-modal, not all of the data are significant enough, especially when the source data vary slowly as compared to data transmission. Therefore, as the first step of semantic signal processing, significance-oriented sampling filters most of unimportant source data, saving communication resources for sequential coding and modulation. On the other hand, even the sampled “significant enough” data may have different level of importance according to the varying environment. Therefore, significance-oriented coding and modulation policy designs provide unequal protection for sampled data with different level of importance. Specifically, more significant data (e.g., status updates with lower age) are given more coding redundancy or more power allocation to ensure their successful acknowledgment at the receiver. By significance-oriented semantic communication system, implementation of semantic communication systems in the SAGSIN (which will be elaborated in the following two sections) will be more energy-efficient.
@@
In this section, we will review another major semantic communication system, called meaning-oriented semantic communication system. Since the primary purpose of such meaning-oriented system design is to achieve perfect recovery of the original data, this meaning-oriented semantic communication system is characterized by a strict symmetrical structure as shown in Fig. 8. In the SAGSIN, even after the semantic-aware sampling mentioned in Section III, there are still a large volume of data due to numerous devices and massive connections. To reduce the transmitting data, it is necessary to perform semantic extraction from the sampled data. Here, the term “semantics” is interpreted as “meaning”. However, the meaning of data is quite implicit to the extent that traditional codes cannot easily capture it. Hopefully, because of the similarity, we may associate meaning extraction with the process of human language learning. Humans gradually grasp semantics through extensive “training” such as teacher instruction, reading books, conversations with others, and so on. Also, a fact is that the original intention of the design for artificial neural networks (ANNs) is to mimic the human brain. Taking the deep neural network (DNN) as an example, through training on large amounts of data, it can learn something from the data to some extent just like humans. Thus, using DNNs to construct the system codec has become the most popular alternative nowadays to “learn” the semantics from original data. Based on above, the meaning-oriented semantic communication system adopting a joint perception-communication-computing framework is established, which is illustrated in Fig. 8. Aiming at conducting intelligent meaning extraction, coding, and modulation, we assume that the input data have been sampled (in a significance-oriented manner). Specifically, DL-based meaning extractor captures the semantic information of sampled data to generate semantic representation, which serves as input of a JSCC encoder and modulator to generate transmitted messages for transmission. At the receiver, after demodulation and JSCC decoding, the semantic representation is recovered, which will be processed by DL-based meaning reconstructor to recover the raw sampled data. During the training of such large-scale neural network, edge servers and cloud data centers must be utilized to provide knowledge bases and facilitate distributed learning, since edge devices are usually limited in computing resources in the SAGSIN scenarios. Also, the knowledge bases distributed among edge/cloud servers can be synchronized via data sharing between transceivers. In the following subsections, we will firstly elaborate on the metrics for representing the system performance in terms of the “meaning similarity” in Section IV-A. Next, in Section IV-B, we will provide a detailed description of the DNNs used to construct the system codec, which strongly intersects with the field of computer science. Finally, we discuss the related techniques for meaning enhancement in Section IV-C.
@@ IV-A Metrics: Towards Meaning Similarity
In this subsection, we review the metrics for the meaning-oriented semantic communication system. A perfect metric can maximally represent the “meaning similarity” between the recovered data at the receiver and the original one at the transmitter. We classify these metrics according to the source type and acquisition method. For the former, we consider text, image/video, and speech sources. As for the latter, based on using whether explicit formulas or neural networks (NNs) to obtain metric values, we analyze objective metrics and learning-based metrics. Please note that the NNs for obtaining metrics is different from the one for building the semantic communication system. The comparison of all the involved metrics is presented in Table III, and these metrics play crucial guiding roles for the system design in Section IV-C.
Sources | Metrics | Explanation | Formula |
@@
Here, the “meaning similarity” can be interpreted as the similar degree of the meaning contained in the texts (which is composed of words and sentences) of the transmitter and receiver. Objective metrics. Similar to the bit error rate (BER) in traditional communication systems, the word error rate (WER) is a fundamental metric to measure the text differences. It is calculated by
@@
where is the number of elements in the -th gram, and is the frequency count function for the -th element in the -th gram. The BLEU score falls within the range of 0 to 1, and a higher score indicates the greater text similarity. Besides, consensus-based image description evaluation (CIDEr)[72] is used to assess the consistency and quality between the reference and generated image descriptions, and thus it can be seen as a metric for evaluating the similarity between texts. Moreover, the metric of semantic similarity (MSS)[73] can measure both the meaning similarity between texts and the correctness of the recovered text (i.e., whether it is a fluent statement), and the contribution weights of the two aspects for the metric can be adjusted. The aforementioned metrics are all based on the text words, and their exploration for text semantics is far from sufficient. Taking the BLEU as an example, sentences “I am available on the weekend.” and “I am not busy during the weekend.” have the same semantics, but the BLEU score for them is not 1. To explore deep semantics of sentences (such as synonyms and expressions with similar meanings), we should resort to learning-based metrics. Learning-based metrics. The bidirectional encoder representations from Transformers (BERT) model[85] is a pre-trained language model that learns rich language representations from a vast amount of text data through large-scale unsupervised training. ††footnotetext: In table III, formulas for some metrics are omitted by * due to space limits in the main text. Readers can go to corresponding initial references for details. The BERT score[74] is a metric built upon the BERT model to evaluate text similarity, utilizing the contextual understanding learned by the BERT neural network. [74] mentions three types of BERT scores, which are recall, precision, and F1 scores222For the definitions of these scores, readers can go to Section V-A for more details.. Considering the need to measure the meaning similarity, the BERT score here specifically refers to the precision score. We assume that the representation vector generated by the embedding model for the original text is denoted as , and the representation vector for the recovered text is denoted as . The BERT score is given by
@@
is between 0 and 1, and a higher score implies the better text similarity. Sentence similarity[75] is another metric based on the BERT model that effectively evaluates the meaning similarity between two sentences. It is given by
where is the BERT model to map the sentence to its semantic vector space. Like the rescaled BERT score, sentence similarity also ranges from 0 to 1. When the similarity between two sentences is maximum, the sentence similarity value is equal to 1. The above two metrics are both based on the BERT model, which leverages the learning from billions of sentences to comprehend certain semantic information within the sentences. Therefore, they can provide a more accurate evaluation of the meaning similarity between texts than objective metrics.
@@ IV-A2 Image (Video) sources
A digital image is composed of a finite number of pixels, while a video is composed of a continuous sequence of images, with each image referred to as a “frame”. Therefore, when discussing metrics, images and videos can be grouped together for analysis, and we take image sources as an example next. The semantics of images heavily rely on context, making it challenging to directly define the “meaning similarity”. For instance, without providing contextual information, we cannot determine whether “a yellow triangle” is more similar to “a green triangle” or “a yellow square”. Objective metrics. Assume that the original and recovered images are respectively and . The matrix dimensions depend on the image properties, such as two dimensions for grayscale images, and three dimensions for colorful images. The peak signal-to-noise ratio (PSNR), comparing each corresponding pixel of two images, is one of the typical metrics to measure the differences between images. The PSNR of a grayscale image is defined as
@@
As for colorful images, one common approach is to calculate the MSE for each color channel separately and then take the average, which is used in (14) to get the PSNR. Obviously, PSNR cannot provide an accurate similarity measurement of colorful images. The structural similarity index (SSIM)[76], considering the similarity of the luminance, the contrast and the structure, is effective in evaluating the colorful image similarity. It is defined as
@@
SSIM is a value between 0 and 1, and a larger value indicates smaller differences between images. The multi-scale structural similarity index (MS-SSIM)[76] is an extension of SSIM. It applies multiple downsampling operations on the original image to obtain images at different scales. Then, the SSIMs for these images are calculated separately, and finally a weighted summation operation of these SSIMs is performed, that is
where , , and are used to adjust the relative importance for different components. MS-SSIM takes into account local details and global structures of images, thereby generally providing a better evaluation of the image quality compared to SSIM. Considering the need for the accurate measurement of meaning similarity between images, the aforementioned three metrics are simple and shallow, and thus cannot capture the subtle differences perceived by humans. Learning-based metrics. The learned perceptual image patch similarity (LPIPS)[77] learns image semantics from the human perception perspective through DNNs. It evaluates image similarity by calculating the distance between the feature representations of images extracted from pretrained DNNs such as the SqueezeNet, AlexNet and VGG. LPIPS is defined as
@@
Smaller values of FID and KID indicate greater meaning similarity between images. They are consistent with human perception and can reveal semantic differences between images at a deeper level. In addition, considering the video conferencing scenario, [80] proposes the average keypoint distance (AKD) based on the keypoints (contain certain semantic information) extracted by a pretrained facial landmark detector to evaluate the differences between images. The learning-based metrics above go beyond the pixel level and instead extract deeper semantic information from images using different pre-trained models. Thus, they can effectively measure the meaning similarity between images.

### v3

@@ III-B3 Significance-oriented modulation policy design
Modulation techniques is crucial in the wireless communication system, for they can enhance the reliability, spectral efficiency, and data transmission rate. However, the modulation technique design encounters challenges in efficiently utilizing limited spectrum resources to ensure timeliness in the SAGSIN. To address this, NOMA has emerged as a promising solution, which enables multiple users to share time-frequency resources by using power domain multiplexing. By adopting power allocation (i.e., assigning varying power levels to different users) which is a modulation policy, NOMA facilitates simultaneous message transmission of multiple users. Allocating more power to the data of higher significance not only ensures reliable transmission of significant data but also enhances the timeliness of the system by directly influencing data rates and system capacity.
@@
Various techniques are employed to enhance the timeliness performance of NOMA systems. [77] considers the problem of energy-efficient scheduling for minimizing the AoI in an opportunistic NOMA/orthogonal-multiple-access (NOMA/OMA) downlink broadcast wireless network. The initial step involves formulating a resource allocation problem to minimize the average AoI in the network. To address energy-efficiency considerations, the work takes into account both a long-term average power constraint and a maximum power constraint. Then the Lyapunov framework is employed to approximate the original problem as a queue stability problem. The obtained single time-slot optimization problem is further decomposed into power allocation and user scheduling subproblems. An efficient piece-wise linear approximation is utilized to solve the non-convex power allocation subproblem, while linear programming algorithm is adopted to solve the user scheduling subproblem. Unlike the approach discussed in [77] that focuses solely on NOMA and OMA transmission schemes, [78] proposes an adaptive NOMA/OMA/cooperative-SWIPT-NOMA transmission scheme. In the NOMA scheme, the base station is responsible for allocating power to each user. However, in the cooperative-SWIPT-NOMA scheme, the base station not only allocates power to each user but also determines the power splitting coefficient (which determines the amount of harvested energy). The authors focus on a short packet communication system within a wireless network, where a base station transmits timely status updates to the users. The objective is to minimize the weighted average AoI by selecting appropriate multiple access techniques and corresponding power allocations. The resource allocation problem for the adaptive transmission scheme is formulated as an MDP and solved by iteration algorithms. Similar to [78], the authors in [79] investigate the AoI performance in a downlink wireless communication system, while HARQ-CC aided NOMA technology is employed in [79]. [79] considers a more flexible scenario where both power allocation decisions and retransmission decisions are made adaptively to minimize the system average AoI in each time slot. An AoI minimization problem is formulated, and the optimal policy is obtained by transforming the problem to MDP. Furthermore, the authors notice the issue of fairness between users, and propose a greedy algorithm to minimize the users’ maximal expected AoI, ensuring fairness by preventing the far user (with weak channel condition) from being deprived of timely service.
The aforementioned works primarily focus on ground-based wireless communication scenarios, while [80, 81] consider the satellite-based IoT (S-IoT) network. In particular, the authors in [80] propose an AoI-minimal resource allocation scheme in NOMA-based S-IoT downlink network. In this system, the satellite sends timely status updates to multiple user equipment devices. To ensure the freshness of the status updates in the network, an AoI optimization problem is formulated, taking into account long-term average power, peak power, and throughput constraints. The problem is then transformed into a series of online power allocation problems using Lyapunov optimization tools. Considering the limited computing resources of satellites, the particle swarm optimization (PSO) algorithm is employed to solve the non-convex optimization problem with linear computational complexity, named NOMA-AoI scheme. In the PSO algorithm, each particle means a possible solution which is the power allocation factors. By iteratively updating the velocities and positions of particles based on their historical information and the globally best information, the PSO algorithm explores the search space to find the optimal power allocation strategy. Based on [80], the authors in [81] further consider the network stability constraint in AoI optimization problem under NOMA-based S-IoT network. The problem is then transformed into three queue stability problems using the Lyapunov framework, which also converts the optimization problem into a series of single time slot deterministic optimization problems. The weights for the queue backlog and channel conditions are derived using the ListNet algorithm, a machine-learning-based approach, to obtain an optimized power allocation order with linear complexity. Similar to [80], the proposed NOMA long-term AoI minimization power allocation problem is also addressed using the PSO algorithm.
@@ III-C Lessons Learned from This Section
This section focuses on significance-oriented sampling, coding, and modulation policies to release heavy burden caused by massive source data and ensure that only significant data are sampled and transmitted for further processing and actuation. In fact, in the SAGSIN scenarios, although data generated by distributed edge devices are massive and multi-modal, not all of the data are significant enough, especially when the source data vary slowly as compared to data transmission. Therefore, as the first step of semantic signal processing, significance-oriented sampling filters most of unimportant source data, saving communication resources for sequential coding and modulation.
Moreover, even the sampled “significant enough” data may have different levels of importance according to the varying environment. Therefore, significance-oriented coding and modulation policy designs provide unequal protection for sampled data with different levels of importance. Specifically, more significant data (e.g., status updates with lower age) are given more coding redundancy or more power allocation to ensure their successful acknowledgment at the receiver. By significance-oriented semantic communication system, implementation of semantic communication systems in the SAGSIN (which will be elaborated in the following two sections) will be more energy-efficient.
Finally, it is also a promising idea to enrich the concept of “significance” metrics such that they are meaning-aware or effectiveness-aware, and study meaning- (effectiveness-) aware communication policy optimization. In fact, there have been pioneering works on refining AoI metric by considering spacial-temporal information correlation[83]. In SAGSIN, such information correlations are usually meaning-related, because the adjacent (in space or time) statuses in real world have similar meanings (such as similar fire statuses among the same area or during a period of time). Also, some preliminary efforts have sought to refine the AoII metric by introducing meaning-oriented metrics (which will be elaborated in the next section) as information penalty function[84]. By such means, the refined AoII can capture the meaning similarity of the statuses between transceivers. Furthermore, if we introduce the cost of decision/actuation as penalty function, effectiveness-aware significance metrics can also be proposed. Note that in cases when we utilize such meaning- (effectiveness-) aware metrics to guide SPT design, we should firstly acquire the meaning (or effectiveness-related information) before sampling/coding/modulation processes. The methods on acquiring meaning and effectiveness-related information will be elaborated on Section IV and Section V, respectively.
@@
In this section, we will review another major semantic communication system, called meaning-oriented semantic communication system. Since the primary purpose of such meaning-oriented system design is to achieve perfect recovery of the original data, this meaning-oriented semantic communication system is characterized by a strict symmetrical structure as shown in Fig. 8.
@@
In the SAGSIN, even after the semantic-aware sampling mentioned in Section III, there are still a large volume of multi-modal source data (such as texts, images, audios, and videos) collected by numerous perception devices. To reduce the amount of transmitting data, it is necessary to perform semantic extraction from the source data. Here, the term “semantics” is interpreted as “meaning”. However, the meaning of data is quite implicit to the extent that traditional codes cannot easily capture it. Hopefully, because of the similarity, we may associate meaning extraction with the process of human language learning. Humans gradually grasp semantics through extensive “training” such as teacher instruction, reading books, conversations with others, and so on. Also, a fact is that the original intention of the design for artificial neural networks (ANNs) is to mimic the human brain. Taking the deep neural network (DNN) as an example, through training on large amounts of data, it can learn something from the data to some extent just like humans. Thus, using DNNs to construct the system codec has become the most popular alternative nowadays to “learn” the semantics from original data.
Based on above, the meaning-oriented semantic communication system adopting a joint perception-communication-computing framework is established, which is illustrated in Fig. 8. Aiming at conducting intelligent semantic extraction, coding, and modulation, we assume that the type of source data are given data in this section and the next section, unless otherwise specified. Specifically, DL-based meaning extractor captures the semantic information of source data to generate semantic representation, which serves as input of a JSCC encoder and modulator to generate transmitted messages for transmission. At the receiver, after demodulation and JSCC decoding, the semantic representation is recovered, which will be processed by DL-based meaning reconstructor to recover the source data. During the training of such large-scale neural networks, edge servers and cloud data centers must be utilized to provide knowledge bases and facilitate distributed learning, since end devices are usually limited in computing resources in the SAGSIN scenarios. Also, the knowledge bases distributed among edge/cloud servers can be synchronized via data sharing between transceivers.
It is worth noting that in distributed learning scenario, the meaning extractor/reconstructor and JSCC encoder/decoder & modulator/demodulator will be firstly pre-trained by various data sets on the edge/cloud servers before being deployed. Secondly, the pre-trained models will be downloaded to the corresponding modules (or devices). Next, according to whether the module parameters are updated during communication processes, different communication modules will adopt either offline learning policy (without parameter updating) or online learning policy (with parameter updating). In online learning policy, module parameters will be also updated based on the distance between real data and the output data, similar to the gradient descent methods in training processes. In the following, unless otherwise specified, the mentioned works on meaning-oriented semantic communications all adopt offline learning methods.
In the following subsections, we will firstly elaborate on the metrics for representing the system performance in terms of the “meaning similarity” in Section IV-A. Next, in Section IV-B, we will provide a detailed description of the DNNs used to construct the system codec, which strongly intersects with the field of computer science. Finally, we discuss the related techniques for meaning enhancement in Section IV-C.
@@ IV-A Metrics: Towards Meaning Similarity
In this subsection, we review the metrics for the meaning-oriented semantic communication system. A perfect metric can maximally represent the “meaning similarity” between the recovered data at the receiver and the original one at the transmitter. We classify these metrics according to the source type and acquisition method. For the former, we consider text, image/video, and speech sources. As for the latter, based on using whether explicit formulas or neural networks (NNs) to obtain metric values, we analyze objective metrics and learning-based metrics. Please note that the NNs for obtaining metrics are different from those for building semantic communication systems. The comparison of all the involved metrics is presented in Table V, and these metrics play crucial guiding roles for the system design in Section IV-C. As will be elaborated later, learning-based metrics are usually more effective in describing the meaning similarity than objective metrics. The reason behind is that the NNs that evaluate the similarity are trained from various aspects, while objective metrics usually demonstrate the similarity from only one of the aspects.
Sources | Metrics | Explanation and Application Scenarios | Formula |
@@ IV-A1 Text sources
Here, the “meaning similarity” can be interpreted as the similar degree of the meaning contained in the texts (which is composed of words and sentences) of the transmitter and receiver.
Objective metrics. Similar to the bit error rate (BER) in traditional communication systems, the word error rate (WER) is a fundamental metric to measure the text differences. It is calculated by
@@
where is the number of elements in the -th gram, and is the frequency count function for the -th element in the -th gram. The BLEU score falls within the range of 0 to 1, and a higher score indicates greater text similarity.
Besides, consensus-based image description evaluation (CIDEr)[86] is used to assess the consistency and quality between the reference and generated image descriptions, and thus it can be seen as a metric for evaluating the similarity between texts. Moreover, the metric of semantic similarity (MSS)[87] can measure both the meaning similarity between texts and the correctness of the recovered text (i.e., whether it is a fluent statement), and the contribution weights of the two aspects for the metric can be adjusted.
The aforementioned metrics are all based on the text words, and their exploration for text semantics is far from sufficient. Taking the BLEU as an example, sentences “I am available on the weekend.” and “I am not busy during the weekend.” have the same semantics, but the BLEU score for them is not 1. To explore deep semantics of sentences (such as synonyms and expressions with similar meanings), we should resort to learning-based metrics.
Learning-based metrics. The bidirectional encoder representations from Transformers (BERT) model[99] is a pre-trained language model that learns rich language representations from a vast amount of text data through large-scale unsupervised training.
The BERT score[88] is a metric built upon the BERT model to evaluate text similarity, utilizing the contextual understanding learned by the BERT neural network. [88] mentions three types of BERT scores, which are recall, precision, and F1 scores222For the definitions of these scores, readers can go to Section V-A for more details.. Considering the need to measure the meaning similarity, the BERT score here specifically refers to the precision score. We assume that the representation vector generated by the embedding model for the original text is denoted as , and the representation vector for the recovered text is denoted as . The BERT score is given by
@@
is between 0 and 1, and a higher score implies the better text similarity.
Sentence similarity[89] is another metric based on the BERT model that effectively evaluates the meaning similarity between two sentences. It is given by
where is the BERT model to map the sentence to its semantic vector space. Like the rescaled BERT score, sentence similarity also ranges from 0 to 1. When the similarity between two sentences is maximum, the sentence similarity value is equal to 1. ††footnotetext: In table V, formulas for some metrics are omitted by * due to space limits in the main text. Readers can go to corresponding initial references for details.
The above two metrics are both based on the BERT model, which leverages the learning from billions of sentences to comprehend certain semantic information within the sentences. Therefore, they can provide a more accurate evaluation of the meaning similarity between texts than objective metrics.
@@ IV-A2 Image (Video) sources
A digital image is composed of a finite number of pixels, while a video is composed of a continuous sequence of images, with each image referred to as a “frame”. Therefore, when discussing metrics, images and videos can be grouped together for analysis, and we take image sources as an example next. The semantics of images heavily rely on context, making it challenging to directly define the “meaning similarity”. For instance, without providing contextual information, we cannot determine whether “a yellow triangle” is more similar to “a green triangle” or “a yellow square”.
Objective metrics. Assume that the original and recovered images are respectively and . The matrix dimensions depend on the image properties, such as two dimensions for grayscale images, and three dimensions for colorful images.
The peak signal-to-noise ratio (PSNR), comparing each corresponding pixel of two images, is one of the typical metrics to measure the differences between images. The PSNR of a grayscale image is defined as
@@
As for colorful images, one common approach is to calculate the MSE for each color channel separately and then take the average, which is used in (14) to get the PSNR. Obviously, PSNR cannot provide an accurate similarity measurement of colorful images.
The structural similarity index (SSIM)[90], considering the similarity of the luminance, the contrast and the structure, is effective in evaluating the colorful image similarity. It is defined as
@@
SSIM is a value between 0 and 1, and a larger value indicates smaller differences between images.
The multi-scale structural similarity index (MS-SSIM)[90] is an extension of SSIM. It applies multiple downsampling operations on the original image to obtain images at different scales. Then, the SSIMs for these images are calculated separately, and finally a weighted summation operation of these SSIMs is performed, that is
where , , and are used to adjust the relative importance for different components. MS-SSIM takes into account local details and global structures of images, thereby generally providing a better evaluation of the image quality compared to SSIM.
Considering the need for the accurate measurement of meaning similarity between images, the aforementioned three metrics are simple and shallow, and thus cannot capture the subtle differences perceived by humans.
Learning-based metrics. The learned perceptual image patch similarity (LPIPS)[91] learns image semantics from the human perception perspective through DNNs. It evaluates image similarity by calculating the distance between the feature representations of images extracted from pretrained DNNs such as the SqueezeNet, AlexNet and VGG. LPIPS is defined as
@@
Smaller values of FID and KID indicate greater meaning similarity between images. They are consistent with human perception and can reveal semantic differences between images at a deeper level.
In addition, considering the video conferencing scenario, [94] proposes the average keypoint distance (AKD) based on the keypoints (containing certain semantic information) extracted by a pretrained facial landmark detector to evaluate the differences between images.
The learning-based metrics above go beyond the pixel level and instead extract deeper semantic information from images using different pre-trained models. Thus, they can effectively measure the meaning similarity between images.